-3(4 - x) = 5 - (x + 1)
Answer by macston(5194)

.
-3(4 - x) = 5 - (x + 1)
-12 + 3x = 5 - x - 1 . Add x to each side.
-12 + 4x = 5-1
-12 + 4x = 4 . Add 12 to each side.
4x = 16 . Divide each side by 4
x = 4